OpenStudy (anonymous):

how do i factor 900-1?

OpenStudy (anonymous):

(900,1) 900 x 1 = 900 (90,10) 90 x 10 = 900 (180,5) etc... (300,3) (450,2) (100,9) (150,6) (75,12) (30,30) (225,4) (25,36) (15,60) (18,50) (45,20)

OpenStudy (anonymous):

Divide by the smallest prime numbers and ascend... 900/2 = 450 450/2 = 225 225/3 = 75 75/3 = 25 25/5 = 5 5/5 = 5 900 = 2²+3²+5² The factors will come in pairs (x*y). For x you can have 2^0, 2 or 2² * 3^0, 3 or 3² * 5^0, 5 or 5² = 3x3x3 = 27. So you can have 27 different factors, and y will not interest us as itr will be a repeat of x's other terms.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

how would i set it up in the (___) (___)

OpenStudy (anonymous):

That I don't know...

OpenStudy (anonymous):

ok but thank you for your help

OpenStudy (anonymous):

Sure.

OpenStudy (mathstudent55):

I don't understand your question. Are you trying to factor 900, or 900-1 which is equal to 899?

OpenStudy (mathstudent55):

If it is 900, see this: https://answers.yahoo.com/question/index?qid=20090915160839AAKoTAM

OpenStudy (anonymous):

thank you